### v2

# Total variation bounds in the Lindeberg central limit theorem

Abstract
In this paper, we obtain an explicit total variation bound in the central limit theorem for the sums of non-i.i.d. random variables. Our results show that, under suitable assumptions, Lindeberg’s condition is sufficient and necessary for the convergence in total variation distance.
Keywords: Central limit theorem, Total variation distance.
2010 Mathematics Subject Classification: 60F15.
1 Introduction
Let $(X_{k})_{k\geq 1}$ be a sequence of independent real-valued random variables (not necessarily identically distributed) with mean $\mathbb{E}X_{k}=0$ and finite variances $\sigma_{k}^{2}=\mathbb{E}X_{k}^{2}\in(0,\infty).$ We consider the normalized sums
$S_{n}=\frac{X_{1}+\cdots+X_{n}}{b_{n}},$ (1.1)
where $b_{n}=\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}+\cdots+\sigma_{n}^{2}},\,\,n\geq 1.$
Let us recall the classical central limit theorem: if the random variables $X_{k}$ satisfies Lindeberg’s condition:
$\lim\limits_{n\to\infty}\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[X_{k}^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]=0$ (1.2)
for all $\varepsilon>0,$ where $\rm 1\hskip-3.87495pt1$} is the indicator function, then $S_{n}$ converges in distribution to a standard normal random variable $N$ as $n\to\infty.$ Furthermore, if the random variables $X_{k}$ satisfy the Feller-Lévy condition
$\max\limits_{1\leq k\leq n}(\sigma_{k}^{2}/b_{n}^{2})\to 0\,\,\,\text{as}\,\,n\to\infty,$ (1.3)
then Lindeberg’s condition is both sufficient and necessary. We also have the following optimal bound on the Kolmogorov distance for the rate of convergence (see [3] for a review)
$d_{K}(S_{n},N):=\sup_{x\in\mathbb{R}}|P(S_{n}\leq x)-P(N\leq x)|$ $\leq\frac{c}{b_{n}^{3}}\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]$ (1.4)
$\leq\frac{c}{b_{n}^{3}}\sum\limits_{k=1}^{n}\mathbb{E}|X_{k}|^{3},$ (1.5)
where $c$ is a positive absolute constant. We now consider the total variation distance between the laws of $S_{n}$ and $N$ defined by
$d_{TV}(S_{n},N):=\frac{1}{2}\sup_{\|h\|_{\infty}\leq 1}|Eh(S_{n})-Eh(N)|,$
where $h$ is measurable and $\|h\|_{\infty}:=\sup\limits_{x\in\mathbb{R}}|h(x)|.$ Note that the convergence in total variation is stronger than the convergence in distribution. The study of the central limit theorem in total variation distance has a long history beginning in 1952 with the results of Prokhorov [12]. However, the problem of obtaining optimal estimates for $d_{TV}(S_{n},N)$ (under minimal moment conditions) is non-trivial. The reader can consults [1, 13] and references therein for the case of independent identically distributed (i.i.d.) random variables with finite third absolute moment. For the case of non-i.i.d. random variables, it is natural to wonder whether the bounds (1.4) and (1.5) may be strengthened to the total variation distance as well.
When all summands $S_{n}$ have discrete distributions, $d_{TV}(S_{n},N)=1\nrightarrow 0$ as $n\to\infty.$ Hence, to obtain an affirmative answer, we need to impose additional assumptions on the sequence $(X_{k})_{k\geq 1}.$ We have the following result, due to Bobkov-Chistyakov-Götze [5]: Assume that the independent random variables $X_{1},\cdots,X_{n}$ have finite third absolute moments, and that $D(X_{k})\leq D$ for all $1\leq k\leq n$ and for some $D>0.$ Then
$d_{TV}(S_{n},N)\leq\frac{C_{D}}{b_{n}^{3}}\sum\limits_{k=1}^{n}\mathbb{E}|X_{k}|^{3},$
where $C_{D}$ is a positive constant depending only on $D.$ Here $D(.)$ denotes the relative entropy defined as follows: for a random variable $X$ with density $p(x)$ and finite variance,
$D(X)=\int_{-\infty}^{\infty}p(x)\log\frac{p(x)}{\varphi_{X}(x)}dx,$ where $\varphi_{X}(x)=\frac{1}{\sqrt{2\pi{\rm Var}(X)}}e^{-\frac{(x-\mathbb{E}X)^{2}}{2{\rm Var}(X)}},x\in\mathbb{R}.$
The aim of this paper is to establish the bound (1.4) for $d_{TV}(S_{n},N).$ Standing for the finiteness of relative entropies, we require the random variables $X_{k}^{\prime}s$ to have the finite standardised Fisher information. We recall that, given a random variable $X$ with an absolutely continuous density $p(x),$ the standardised Fisher information of $X$ (or its distribution) is defined by
$J_{st}(X)={\rm Var}(X)\int_{-\infty}^{+\infty}\frac{p^{\prime}(x)^{2}}{p(x)}dx-1={\rm Var}(X)\mathbb{E}[\rho^{2}(X)]-1,$
where $p^{\prime}$ denotes a Radon-Nikodym derivative of $p$ and $\rho:=p^{\prime}/p$ is the score function. For more details about the Fisher information and the central limit theorem results, we refer the reader to the monograph [9] and recent papers [6, 8, 10].
Hereafter, we put $J(X)=J_{st}(X)+1.$ Our main results are stated in the following theorem.
Theorem 1.1.
Let $(X_{k})_{k\geq 1}$ be a sequence of independent real-valued random variables $\mathbb{E}X_{k}=0$ and finite variances $\sigma_{k}^{2}=\mathbb{E}X_{k}^{2}\in(0,\infty).$ Assume that, for every $k\geq 1,$ the law of $X_{k}$ has an absolutely continuous density $p_{k}$ satisfying $J(X_{k})<\infty.$ Then, for $S_{n}$ defined by (1.1), we have
$d_{TV}(S_{n},N)\leq\bigg(\frac{8\pi\max\limits_{1\leq k\leq n}J(X_{k})}{1-\max\limits_{1\leq k\leq n}(\sigma_{k}^{2}/b_{n}^{2})}\bigg)^{\frac{1}{2}}\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{n}^{3}}.$ (1.6)
Our condition $J(X_{k})<\infty$ is stronger than the condition $D(X_{k})<\infty$ required in [5]. Indeed, we have $D(X_{k})\leq(J(X_{k})-1)/2,$ see e.g. Lemma E.3 in [9]. However, our bound (1.6) allows us to upgrade the Lindeberg central limit theorem to the convergence in total variation distance as follows.
Corollary 1.1.
Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem 1.1. Assume additionally that $\sup\limits_{k\geq 1}J(X_{k})<\infty.$ The, the following two statements are equivalent
$(i)$ Lindeberg’s condition (1.2) is satisfied.
$(ii)$ $d_{TV}(S_{n},N)\to 0$ as $n\to\infty$ and the Feller-Lévy condition (1.3) is satisfied.
We use the condition $J(X_{k})<\infty$ is to establish the bound (1.6). However, this condition can be removed to recover the classical central limit theorem.
Corollary 1.2.
Let $(Y_{k})_{k\geq 1}$ be a sequence of independent real-valued random variables $\mathbb{E}Y_{k}=0$ and finite variances $\alpha_{k}^{2}=\mathbb{E}Y_{k}^{2}\in(0,\infty).$ Assume that this sequence satisfies Lindeberg’s condition:
$\lim\limits_{n\to\infty}\frac{1}{a_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[Y_{k}^{2}{\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|>\varepsilon a_{n}\}}]=0$ (1.7)
for all $\varepsilon>0,$ where $a_{n}:=\sqrt{\alpha_{1}^{2}+\cdots+\alpha_{n}^{2}}.$ Then, as $n\to\infty,$ we have $\frac{Y_{1}+\cdots+Y_{n}}{a_{n}}\to N$ in distribution.
2 Proofs
We first prepare some technical results. For a given real valued measurable function $h$ with $\mathbb{E}h(Z)<\infty.$ The ordinary differential equation
$f^{\prime}(x)-xf(x)=h(x)-\mathbb{E}h(Z)$ (2.1)
is called Stein’s equation associated with $h$. The unique bounded solution of (2.1) is given by
$f(x):=e^{x^{2}/2}\int_{-\infty}^{x}(h(y)-\mathbb{E}h(Z))e^{-y^{2}/2}dy,\,\,x\in\mathbb{R}.$ (2.2)
We recall from Lemma 2.4 in [7] the following estimates for $f.$
Lemma 2.1.
Let $f$ be the solution (2.2) to Stein’s equation. If $h$ is bounded $1,$ then
$\text{$\|f\|_{\infty}\leq\sqrt{2\pi}$ and $\|f^{\prime}\|_{\infty}\leq 4$}.$
Lemma 2.2.
Let the sequence $(X_{k})_{k\geq 1}$ be as in Theorem 1.1. For each $1\leq k\leq n,$ we put
$S_{k,n}=\frac{X_{1}+\cdots+X_{k-1}+X_{k+1}+\cdots+X_{n}}{b_{k,n}},$
where $b_{k,n}=\sqrt{b_{n}^{2}-\sigma_{k}^{2}}.$ Then the law of $S_{k,n}$ has an absolutely continuous density. Moreover, the score function $\rho_{k,n}$ of $S_{k,n}$ is given by
$\rho_{k,n}(x)=\mathbb{E}\left[\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})}{b_{k,n}}\bigg|S_{k,n}=x\right],$ (2.3)
where $\rho_{i}$ denotes the score function of $X_{i}.$
Proof.
For each $k\geq 1,$ it follows from Theorem VIII.1 in [2] that the density $p_{k}$ of the random variable $X_{k}$ satisfies
$\lim\limits_{x\to\pm\infty}(1+|x|)p_{k}(x)=0.$
Hence, for any bounded and differentiable function $f:\mathbb{R}\to\mathbb{R}$ with bounded derivative, we can use Lemma 1.20 in [9] to get
$\mathbb{E}[f(X_{k})\rho_{k}(X_{k})]=-\mathbb{E}[f^{\prime}(X_{k})].$
Consequently, by the independence of $X_{k}^{\prime}s,$ we obtain
$\mathbb{E}\left[f(S_{k,n})\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})}{b_{k,n}}\right]$ $=\frac{\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\mathbb{E}\left[f(S_{k,n})\rho_{i}(X_{i})\right]}{b_{k,n}}$
$=-\mathbb{E}[f^{\prime}(S_{k,n})].$ (2.4)
Then, by using the same arguments as in the proof of Proposition 3.2 in our recent paper [8], we conclude that the law of $S_{k,n}$ has an absolutely continuous density and its score function is given by (2.3). This finishes the proof of the lemma. ∎
Definition 2.1.
Given a random variable $Y\in L^{1}(\Omega)$ with a density $p_{Y}$ supported on some connected interval of $\mathbb{R},$ we define the Stein kernel $\tau_{Y}$ associated to $Y$ as follows
$\tau_{Y}(x)=\frac{\int_{x}^{\infty}(y-\mathbb{E}[Y])p_{Y}(y)dy}{p_{Y}(x)}$
for $x$ in the support of $p_{Y}.$
In the proofs, we use the following fundamental properties of Stein kernels, see e.g. [11] and references therein: $\tau_{Y}(Y)\geq 0,$ $\mathbb{E}[\tau_{Y}(Y)]={\rm Var}(Y)$ and
$\mathbb{E}[f(Y)(Y-\mathbb{E}Y)]=\mathbb{E}[f^{\prime}(Y)\tau_{Y}(Y)]$ (2.5)
for any differentiable function $f:\mathbb{R}\to\mathbb{R}$ with bounded derivative.
We now are ready to prove the main results of the present paper.
Proof of Theorem 1.1. We will carry out the proof in two steps.
Step 1. In this step, we additionally assume that
For every $k\geq 1,$ the density $p_{k}$ of $X_{k}$ is supported on some connected interval of $\mathbb{R}.$ (2.6)
According to Definition 2.1, this assumption allows us to define the Stein kernels
$\tau_{k}(X_{k})=\frac{\int_{X_{k}}^{\infty}yp_{k}(y)dy}{p_{k}(x)},\,\,k\geq 1.$
Let $h$ be a measurable function bounded by $1.$ From the Stein equation (2.1) we have
$\mathbb{E}h(S_{n})-\mathbb{E}h(N)$ $=\mathbb{E}f^{\prime}(S_{n})-\mathbb{E}[S_{n}f(S_{n})]$
$=\mathbb{E}f^{\prime}(S_{n})-\frac{1}{b_{n}}\sum\limits_{k=1}^{n}\mathbb{E}[X_{k}f(S_{n})].$
Hence, by the property (2.5) of the Stein kernels, we deduce
$\mathbb{E}h(S_{n})-\mathbb{E}h(N)$ $=\mathbb{E}f^{\prime}(S_{n})-\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[\tau_{k}(X_{k})f^{\prime}(S_{n})]$
$=\frac{1}{b_{n}^{2}}\mathbb{E}\left[f^{\prime}(S_{n})(b_{n}^{2}-\tau_{1}(X_{1})-\cdots-\tau_{n}(X_{n}))\right]$
$=\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right].$ (2.7)
Fixed $1\leq k\leq n.$ Let $S_{k,n}$ be as in Lemma 2.2. By the independence of $X_{k}$ and $S_{k,n},$ we can use the relation (2.4) to get
$\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]$ $=\mathbb{E}\left[f^{\prime}\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]$
$=-\frac{b_{n}}{b_{k,n}}\mathbb{E}\left[f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)(\sigma_{k}^{2}-\tau_{k}(X_{k}))\rho_{k,n}(S_{k,n})\right].$
On the other hand, since $\mathbb{E}[\tau_{k}(X_{k})]=\sigma_{k}^{2},$ we have
$\mathbb{E}\left[f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)(\sigma_{k}^{2}-\tau_{k}(X_{k}))\rho_{k,n}(S_{k,n})\right]=\mathbb{E}\left[f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\rho_{k,n}(S_{k,n})\right]\mathbb{E}[\sigma_{k}^{2}-\tau_{k}(X_{k})]=0.$
So we obtain
$\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]$
$=-\frac{b_{n}}{b_{k,n}}\mathbb{E}\left[\bigg(f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg)(\sigma_{k}^{2}-\tau_{k}(X_{k}))\rho_{k,n}(S_{k,n})\right].$ (2.8)
Furthermore, we have
$\bigg|f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg|$
$=\bigg|f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg|{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}+\bigg|f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg|{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq b_{n}\}}.$
Then, by using the estimates $\|f\|_{\infty}\leq\sqrt{2\pi}$ and $\|f^{\prime}\|_{\infty}\leq 4$ from Lemma 2.1, we deduce
$\bigg|f\bigg(\frac{b_{k,n}S_{k,n}+X_{k}}{b_{n}}\bigg)-f\bigg(\frac{b_{k,n}S_{k,n}}{b_{n}}\bigg)\bigg|$ $\leq\sqrt{8\pi}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}+\frac{4|X_{k}|}{b_{n}}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq b_{n}\}}$
$\leq\frac{\sqrt{8\pi}}{b_{n}}(b_{n}\wedge|X_{k}|).$ (2.9)
Inserting (2.9) into (2.8) yields
$|\mathbb{E}$ $\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]|\leq\frac{\sqrt{8\pi}}{b_{k,n}}\mathbb{E}\left[(b_{n}\wedge|X_{k}|)|(\sigma_{k}^{2}-\tau_{k}(X_{k}))\rho_{k,n}(S_{k,n})|\right]$
$\leq\frac{\sqrt{8\pi}}{b_{k,n}}\mathbb{E}\left[(b_{n}\wedge|X_{k}|)|\sigma_{k}^{2}-\tau_{k}(X_{k})|\right]\mathbb{E}|\rho_{k,n}(S_{k,n})|$
$\leq\frac{\sqrt{8\pi}}{b_{k,n}}\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}\left[(b_{n}\wedge|X_{k}|)\tau_{k}(X_{k})\right]+\frac{\sqrt{8\pi}}{b_{k,n}}\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}\left[(b_{n}\wedge|X_{k}|)\sigma_{k}^{2}\right].$ (2.10)
Once again, we use the property (2.5) of the Stein kernels to get the following estimate for the first addend in the right hand side of (2.10)
$\mathbb{E}[(b_{n}\wedge|X_{k}|)\tau_{k}(X_{k})]$ $=\mathbb{E}\left[X_{k}\int_{0}^{X_{k}}(b_{n}\wedge|y|)dy\right]$
$\leq\mathbb{E}\left[|X_{k}|\int_{0}^{|X_{k}|}(b_{n}\wedge y)dy\right]$
$\leq\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)].$
For the second addend in the right hand side of (2.10), by Chebyshev’s association inequality, we have
$\mathbb{E}[(b_{n}\wedge|X_{k}|)\sigma_{k}^{2}]=\mathbb{E}[(b_{n}\wedge|X_{k}|)]\mathbb{E}[|X_{k}|^{2}]\leq\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)].$
So it holds that
$|\mathbb{E}\left[f^{\prime}(S_{n})(\sigma_{k}^{2}-\tau_{k}(X_{k}))\right]|\leq\frac{2\sqrt{8\pi}}{b_{k,n}}\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)],\,\,1\leq k\leq n.$
From the above estimates and (2.7), we conclude that
$|\mathbb{E}h(S_{n})-\mathbb{E}h(N)|\leq\frac{2\sqrt{8\pi}}{b_{n}^{2}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{k,n}}$
for any measurable function $h$ bounded by $1.$
Step 2. In this step, we remove the assumption (2.6) and conclude the proof. Let $(N_{k})_{k\geq 1}$ be a sequence of independent standard normal random variables, and independence of $(X_{k})_{k\geq 1}.$ Fixed $\delta>0,$ we put $\tilde{X}_{k}=X_{k}+\delta N_{k}$ and
$\tilde{S}_{n}=\frac{\tilde{X}_{1}+\cdots+\tilde{X}_{n}}{\tilde{b}_{n}},\,\,\,\tilde{S}_{k,n}=\frac{\tilde{X}_{1}+\cdots+\tilde{X}_{k-1}+\tilde{X}_{k+1}+\cdots+\tilde{X}_{n}}{\tilde{b}_{k,n}},$
where $\tilde{b}_{n}=\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}+\cdots+\sigma_{n}^{2}+n\delta^{2}}$ and $\tilde{b}_{k,n}=\sqrt{\tilde{b}_{n}^{2}-\sigma_{k}^{2}-\delta^{2}}.$ Since the sequence $(\tilde{X}_{k})_{k\geq 1}$ fulfils the assumption (2.6), this allows to use the result proved in Step 1 and we obtain
$|\mathbb{E}h(\tilde{S}_{n})-\mathbb{E}h(N)|\leq\frac{2\sqrt{8\pi}}{\tilde{b}_{n}^{2}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}|\tilde{\rho}_{k,n}(\tilde{S}_{k,n})|\mathbb{E}[|\tilde{X}_{k}|^{2}(\tilde{b}_{n}\wedge|\tilde{X}_{k}|)]}{\tilde{b}_{k,n}}$ (2.11)
for any measurable function $h$ bounded by $1,$ where $\tilde{\rho}_{k,n}$ denotes the score function of $\tilde{S}_{k,n}.$
Obviously, we have $\lim\limits_{\delta\to 0}\tilde{b}_{n}^{2}=b_{n}^{2},$ $\lim\limits_{\delta\to 0}\tilde{b}_{k,n}=b_{k,n}$ and $\lim\limits_{\delta\to 0}\mathbb{E}[|\tilde{X}_{k}|^{2}(\tilde{b}_{n}\wedge|\tilde{X}_{k}|)]=\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)].$ We also have $\lim\limits_{\delta\to 0}\mathbb{E}|\tilde{\rho}_{k,n}(\tilde{S}_{k,n})|)=\mathbb{E}|\rho_{k,n}(S_{k,n})|$ by Theorem 1.4 in [4]. As a consequence, by letting $\delta\to 0,$ we obtain from (2.11) that
$|\mathbb{E}h(S_{n})-\mathbb{E}h(N)|\leq\frac{2\sqrt{8\pi}}{b_{n}^{2}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}|\rho_{k,n}(S_{k,n})|\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{k,n}}$
for any continuous function $h$ bounded by $1.$ In addition, from the representation formula (2.3), we have
$\mathbb{E}|\rho_{k,n}(S_{k,n})|$ $\leq\sqrt{\mathbb{E}|\rho_{k,n}(S_{k,n})|^{2}}\leq\frac{1}{b_{k,n}}\bigg(\mathbb{E}\bigg|\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}\rho_{i}(X_{i})\bigg|^{2}\bigg)^{\frac{1}{2}}$
$=\frac{1}{b_{k,n}}\bigg(\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{4}\mathbb{E}|\rho_{i}(X_{i})|^{2}\bigg)^{\frac{1}{2}}=\frac{1}{b_{k,n}}\bigg(\sum\limits_{1\leq i\leq n,i\neq k}\sigma_{i}^{2}J(X_{i})\bigg)^{\frac{1}{2}}$
$\leq\big(\max\limits_{1\leq k\leq n}J(X_{k})\big)^{\frac{1}{2}}.$
We therefore obtain, for any continuous function $h$ bounded by $1,$
$|\mathbb{E}h(S_{n})-\mathbb{E}h(N)|$ $\leq\frac{2\big(8\pi\max\limits_{1\leq k\leq n}J(X_{k})\big)^{\frac{1}{2}}}{b_{n}^{2}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{k,n}}$
$=\frac{2\big(8\pi\max\limits_{1\leq k\leq n}J(X_{k})\big)^{\frac{1}{2}}}{b_{n}^{3}}\sum\limits_{k=1}^{n}\frac{\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{\big(1-\sigma_{k}^{2}/b_{n}\big)^{\frac{1}{2}}}$
$\leq 2\bigg(\frac{8\pi\max\limits_{1\leq k\leq n}J(X_{k})}{1-\max\limits_{1\leq k\leq n}(\sigma_{k}^{2}/b_{n}^{2})}\bigg)^{\frac{1}{2}}\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{n}^{3}}.$ (2.12)
By an application of Lusin’s theorem,
$d_{TV}(S_{n},N)=\frac{1}{2}\sup_{h\in\mathcal{C},\|h\|_{\infty}\leq 1}|Eh(S_{n})-Eh(N)|,$
where $\mathcal{C}$ denotes the space of continuous functions. So the bound (1.6) follows directly from (2.12). The proof of Theorem 1.1 is complete. $\square$
Proof of Corollary 1.1. The implication $(ii)\Rightarrow(i)$ is obvious because of the fact that the convergence in total variation implies the convergence in distribution. To verify the implication $(i)\Rightarrow(ii),$ we observe that, for any $\varepsilon\in(0,1),$ we have
$\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}(b_{n}\wedge|X_{k}|)]}{b_{n}^{3}}=\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}]}{b_{n}^{2}}+\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{3}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq b_{n}\}}]}{b_{n}^{3}}$
$=\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}]}{b_{n}^{2}}+\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{3}{\rm 1\hskip-3.87495pt1}_{\{\varepsilon b_{n}<|X_{k}|\leq b_{n}\}}]}{b_{n}^{3}}+\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{3}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq\varepsilon b_{n}\}}]}{b_{n}^{3}}$
$\leq\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>b_{n}\}}]}{b_{n}^{2}}+\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{\varepsilon b_{n}<|X_{k}|\leq b_{n}\}}]}{b_{n}^{2}}+\frac{\varepsilon\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{3}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|\leq\varepsilon b_{n}\}}]}{b_{n}^{2}}$
$\leq\frac{\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]}{b_{n}^{2}}+\varepsilon.$
On the other hand, it is well known that Lindeberg’s condition implies the Feller-Lévy condition (1.3). Hence, recalling the bound (1.6), we deduce
$\lim\limits_{n\to\infty}d_{TV}(S_{n},N)\leq\big(8\pi\sup\limits_{k\geq 1}J(X_{k})\big)^{\frac{1}{2}}\varepsilon.$
Letting $\varepsilon\to 0$ gives us the desired conclusion. $\square$
Proof of Corollary 1.2. Let $(N_{k})_{k\geq 1}$ be a sequence of independent normal random variables, and independence of $(Y_{k})_{k\geq 1}.$ For each $k\geq 1,$ $\mathbb{E}N_{k}=0$ and $\mathbb{E}N_{k}^{2}=\mathbb{E}Y_{k}^{2}.$ We consider the sequence $X_{k}=Y_{k}+N_{k},k\geq 1$ and define $S_{n}$ as in (1.1). For every $\varepsilon>0,$ we have
$\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]$ $=\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|Y_{k}+N_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|Y_{k}+N_{k}|>\varepsilon b_{n}\}}]$
$\leq\frac{2}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[(|Y_{k}|^{2}+|N_{k}|^{2}){\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|+|N_{k}|>\varepsilon b_{n}\}}]$
$\leq\frac{2}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[(|Y_{k}|^{2}+|N_{k}|^{2})({\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|>\varepsilon b_{n}/2\}}+{\rm 1\hskip-3.87495pt1}_{\{|N_{k}|>\varepsilon b_{n}/2\}})].$
By Chebyshev’s association inequality we deduce
$\frac{1}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|X_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|X_{k}|>\varepsilon b_{n}\}}]$ $\leq\frac{4}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|Y_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|Y_{k}|>\varepsilon b_{n}/2\}}]+\frac{4}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|N_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|N_{k}|>\varepsilon b_{n}/2\}}].$
It is easy to check that $\frac{4}{b_{n}^{2}}\sum\limits_{k=1}^{n}\mathbb{E}[|N_{k}|^{2}{\rm 1\hskip-3.87495pt1}_{\{|N_{k}|>\varepsilon b_{n}/2\}}]\to 0$ as $n\to\infty.$ Hence, by the condition (1.7), we conclude that the sequence $(X_{k})_{k\geq 1}$ also satisfies Lindeberg’s condition.
On the other hand, by Lemma 1.20 in [9], the score function of $X_{k}$ is given by
$\rho_{k}(X_{k})=\mathbb{E}\left[\frac{-N_{k}}{\alpha_{k}^{2}}\bigg|X_{k}\right],\,\,k\geq 1.$
This implies that $\sup\limits_{k\geq 1}J(X_{k})<\infty.$ Indeed, we have
$J(X_{k})={\rm Var}(X_{k})\mathbb{E}[\rho_{k}^{2}(X_{k})]=2\alpha_{k}^{2}\mathbb{E}\left(\mathbb{E}\left[\frac{-N_{k}}{\alpha_{k}^{2}}\bigg|X_{k}\right]\right)^{2}\leq 2,\,\,k\geq 1.$
Thanks to Corollary 1.1, we obtain $d_{TV}(S_{n},N)\to 0$ as $n\to\infty.$ For every $t\in\mathbb{R},$ in view of Lévy’s continuity theorem, we have
$\mathbb{E}\left[\exp\left(it\frac{Y_{1}+\cdots+Y_{n}}{a_{n}}\right)\right]$ $=e^{t^{2}/2}\mathbb{E}\left[\exp\left(it\frac{Y_{1}+N_{1}+\cdots+Y_{n}+N_{n}}{a_{n}}\right)\right]$
$=e^{t^{2}/2}\mathbb{E}\left[\exp\left(it\sqrt{2}S_{n}\right)\right]\to e^{-t^{2}/2},\,\,n\to\infty$
and hence, $\frac{Y_{1}+\cdots+Y_{n}}{a_{n}}\to N$ in distribution.
The proof is complete. $\square$